# Beyond Logit Lens: Contextual Embeddings for Robust Hallucination Detection & Grounding in VLMs

### v1

@@ Abstract
The rapid development of Large Multimodal Models (LMMs) has significantly advanced multimodal understanding by harnessing the language abilities of Large Language Models (LLMs) and integrating modality-specific encoders. However, LMMs are plagued by hallucinations that limit their reliability and adoption. While traditional methods to detect and mitigate these hallucinations often involve costly training or rely heavily on external models, recent approaches utilizing internal model features present a promising alternative. In this paper, we critically assess the limitations of the state-of-the-art training-free technique, the logit lens, in handling generalized visual hallucinations. We introduce a refined method that leverages contextual token embeddings from middle layers of LMMs. This approach significantly improves hallucination detection and grounding across diverse categories, including actions and OCR, while also excelling in tasks requiring contextual understanding, such as spatial relations and attribute comparison. Our novel grounding technique yields highly precise bounding boxes, facilitating a transition from Zero-Shot Object Segmentation to Grounded Visual Question Answering. Our contributions pave the way for more reliable and interpretable multimodal models.
@@
Anirudh Phukan3, Divyansh1, Harshit Kumar Morj2, Vaishnavi1, Apoorv Saxena3, Koustava Goswami3, 1IIT Kanpur, 2IIT Bombay, 3Adobe Research Correspondence: phukan@adobe.com
@@ 1 Introduction
Figure 1: Visual Hallucination Detection. Logit lens, which only verifies the presence of the token "Green" in the image, mistakenly considers the answer correct. In contrast, our method assigns a low score to "Green" as it correctly contextualizes the color to the jacket rather than the hair.
@@
Detecting and mitigating hallucinations have been extensively explored in both language Tonmoy et al. (2024) and vision domains Liu et al. (2024a). In the context of language, approaches for providing evidence to support model outputs are often framed as attribution or citation Gao et al. (2023); Huo et al. (2023). In computer vision, similar methods are typically studied under Grounded Visual Question Answering Zhang et al. (2024); Khoshsirat and Kambhamettu (2023). However, many existing techniques necessitate either training from scratch or fine-tuning Jiang et al. (2024), and frequently rely on external models such as retrievers or object detectors. These requirements impose significant training costs or result in increased latency during inference, posing challenges to their practical deployment in real-world scenarios.
Recent studies have increasingly leveraged internal model features, such as latent representations, logits, and attention weights, to tackle the objectives of detecting hallucinations Azaria and Mitchell (2023); Varshney et al. (2023) and grounding generated outputs in LLMs Phukan et al. (2024); Qi et al. (2024). Notably, one such study demonstrates the efficacy of a training-free interpretability technique, known as the logit lens, in identifying and mitigating object hallucinations within Visual-Language Models (VLMs) Jiang et al. (2024).
However, we identify several fundamental limitations with the application of the logit lens in this context, which hinder its effectiveness in addressing more general forms of visual hallucinations as categorized by Yan et al. (2024). Specifically, the authors employ the logit lens nostalgebraist (2020) to probe each image patch to determine the presence of objects. We contend that the reliance on token embeddings from the unembedding layer—which are neither contextual nor easily combinable to form multi-token concepts—results in its failure to handle more generalized hallucination scenarios. Consequently, we propose leveraging token embeddings from the middle layers, which have been shown to be contextual Phukan et al. (2024) and effective in representing concepts Wendler et al. (2024).
@@
  • We extend the method by replacing the use of the logit lens with contextual embeddings from the middle layers of the VLM, leading to successful hallucination detection in categories that previously performed near random and improving detection accuracy in other categories.
@@ 2.2 Attribution and Grounded Visual Question Answering
Our approach is inspired by the training-free object hallucination detection and mitigation method proposed by Jiang et al. (2024), but with a crucial distinction. We utilize output token embeddings from intermediate layers instead of employing the logit lens probe. This modification enables us to overcome several limitations of the previous work, particularly by enhancing the representation of multi-token concepts and contextual relationships (§4.2).
@@ 4.2.1 Hallucination Detection
Given the answer tokens generated by the model, we compute the average embedding at a specific layer $l_{T}$, denoted as $h_{\text{Ans}}$. Formally, let $\{t_{1},t_{2},\ldots,t_{k}\}$ be the sequence of answer tokens, and $h_{l_{T}}(t_{i})$ be the embedding of token $t_{i}$ at layer $l_{T}$. The average embedding $h_{\text{Ans}}$ is computed as:
$h_{\text{Ans}}=\frac{1}{k}\sum_{i=1}^{k}h_{l_{T}}(t_{i})$
Next, we evaluate the similarity between the answer embedding $h_{\text{Ans}}$ and each image patch embedding from a specific layer $l_{I}$. Let $\{p_{1},p_{2},\ldots,p_{n}\}$ denote the set of image patches, and $h_{l_{I}}(p_{j})$ be the embedding of patch $p_{j}$ at layer $l_{I}$. We compute the score for each patch $p_{j}$ as the cosine similarity:
$\text{Score}(p_{j})=\text{CosineSim}(h_{\text{Ans}},h_{l_{I}}(p_{j}))$
@@ 4.2.2 Grounded Visual Question Answering
Random LL Out Probs Ours Random LL Out Probs Ours
@@
Table 1: Comparison of mAP scores across different hallucination categories in the HQH dataset.
@@ 6.1 Hallucination Detection
The performance of our method (§4.2.1) and the baseline methods (§5.3.1) on the HQH dataset is summarized in Table 1. Each category comprises 500 image-question pairs. To select the best layer combination for our method, we created a test and validation split with 250 examples each. The scores in Table 1 are reported for the test split.
@@
Our method, which employs contextual embeddings, fares significantly better in these categories compared to the logit lens approach and often outperforms even output probabilities, sometimes by a substantial margin.
For categories such as Action and OCR, where grounding actions (e.g., "Looking at the Television") or text (e.g., "Melrose Ave.") is necessary, the logit lens technique performs better than random but still falls short compared to output probabilities. This is likely due to the logit lens’s poor handling of multi-token objects. While objects which are usually single token may align relatively well, the complexity of representing actions or texts across multiple tokens poses challenges. Our approach, leveraging middle-layer embeddings known to better represent multi-token concepts Wendler et al. (2024), consistently outperforms the logit lens and, in many instances, surpasses output probabilities.
Interestingly, output probabilities excel in the Count category compared to both our method and the logit lens. This category involves questions about the number of instances of a specific object. The inherent challenge for our approach and the logit lens is that visual elements often have low semantic overlap with numerical tokens (e.g., a "car" in the image is unlikely to semantically match "two" in the text). Consequently, output probabilities provide more accurate hallucination detection in such scenarios.
@@ 6.2.1 Qualitative Examples
Some qualitative results of our bounding box visual grounding technique are showcased in Figure 3. For instance, in Figure 3(a), we ground the phrase "Big Ben" to its corresponding location in the image. Our method effectively highlights the relevant portion by leveraging the latent conceptual representation of "Big Ben" intrinsic to the LMM, facilitating the mapping between textual descriptions and visual concepts. Another example is depicted in Figure 3(b). Remarkably, our method proficiently grounds answers within charts and infographics, a novel capability within the community to the best of our knowledge, achievable in a training-free manner. Additional qualitative examples are available in Appendix C.
@@ 7 Discussion & Conclusion
In this paper, we presented a novel approach for detecting hallucinations in LMMs by leveraging contextual token embeddings from intermediate layers. We identified limitations in the state-of-the-art training-free technique, the logit lens, notably its poor handling of complex visual hallucinations involving attributes, comparisons, and spatial relations due to its reliance on non-contextual embeddings. Our proposed method employs contextual embeddings known to capture richer semantic information, improving hallucination detection across diverse categories.
Our experimental results on the HQH benchmark showed that our method enables detection in categories that previously performed near random and improves detection for multi-token concept representation. Furthermore, by introducing a novel grounding technique with highly precise bounding boxes, we advanced Zero-Shot Object Segmentation to a more general GVQA task, validating our approach’s effectiveness on TextVQA-X and VizWiz-G datasets.
@@ 8 Limitations & Future Work
Firstly, the validation of both hallucination detection and grounding has been conducted primarily on factual short answer VQA datasets. Extending this to more diverse and complex datasets remains an area for exploration. Secondly, our method is outperformed by output probabilities in the Count category. Future work could explore extensions to the method for abstractive scenarios such as counting. Lastly, although our method excels at highlighting precise evidence regions sufficient for human verification, improving recall remains an open challenge. Subsequent research could investigate leveraging our scoring mechanisms as priors for pre-trained segmentation models or improving the evaluation metrics for grounded visual question answering.
@@ B.2 Grounded Visual Question Answering
Appendix C Additional Qualitative examples for Grounded Visual Question Answering

### v2

@@ Abstract
The rapid development of Large Multimodal Models (LMMs) has significantly advanced multimodal understanding by harnessing the language abilities of Large Language Models (LLMs) and integrating modality-specific encoders. However, LMMs are plagued by hallucinations that limit their reliability and adoption. While traditional methods to detect and mitigate these hallucinations often involve costly training or rely heavily on external models, recent approaches utilizing internal model features present a promising alternative. In this paper, we critically assess the limitations of the state-of-the-art training-free technique, the logit lens, in handling generalized visual hallucinations. We introduce ContextualLens, a refined method that leverages contextual token embeddings from middle layers of LMMs. This approach significantly improves hallucination detection and grounding across diverse categories, including actions and OCR, while also excelling in tasks requiring contextual understanding, such as spatial relations and attribute comparison. Our novel grounding technique yields highly precise bounding boxes, facilitating a transition from Zero-Shot Object Segmentation to Grounded Visual Question Answering. Our contributions pave the way for more reliable and interpretable multimodal models.
@@
Anirudh Phukan3, Divyansh1*, Harshit Kumar Morj2*, Vaishnavi1*, Apoorv Saxena3, Koustava Goswami3, 1IIT Kanpur, 2IIT Bombay, 3Adobe Research Correspondence: phukan@adobe.com
footnotetext: Authors marked with * were interns at Adobe Research when this work was done.
@@ 1 Introduction
Figure 1: Visual Hallucination Detection. Logit lens, which only verifies the presence of the token "Green" in the image, mistakenly considers the answer correct. In contrast, ContextualLens assigns a low score to "Green" as it correctly contextualizes the color to the jacket rather than the hair.
@@
Detecting and mitigating hallucinations has been extensively explored in both language Tonmoy et al. (2024) and vision domains Liu et al. (2024a). In the context of language, attribution and citation have been proposed as approaches for providing evidence to support model output Gao et al. (2023); Huo et al. (2023). In computer vision, similar methods are typically studied under Grounded Visual Question Answering Zhang et al. (2024); Khoshsirat and Kambhamettu (2023). However, many existing techniques necessitate either training from scratch or fine-tuning Jiang et al. (2024), and frequently rely on external models such as retrievers or object detectors. These requirements impose significant training costs or result in increased latency during inference, posing challenges to their practical deployment in real-world scenarios.
Recent studies have increasingly leveraged internal model features, such as latent representations, logits, and attention weights, to tackle the objectives of detecting hallucinations Azaria and Mitchell (2023); Varshney et al. (2023) and ensuring the grounding of generated outputs in LLMs Phukan et al. (2024); Qi et al. (2024). Notably, one such study demonstrates the efficacy of a training-free interpretability technique, known as the logit lens, in identifying and mitigating object hallucinations within Visual-Language Models (VLMs) Jiang et al. (2024).
The logit lens technique, originally introduced in the context of language models, involves directly mapping intermediate activations to the vocabulary space using the unembedding layer, allowing for an interpretable view of token predictions at different layers nostalgebraist (2020). By applying this technique to VLMs, the authors probe each image patch to determine the presence of objects.
However, we identify several fundamental limitations with the application of the logit lens in this context, which hinder its effectiveness in addressing more general forms of visual hallucinations as categorized by Yan et al. (2024). We contend that the reliance on token embeddings from the unembedding layer—which are neither contextual nor easily combinable to form multi-token concepts—results in its failure to handle more generalized hallucination scenarios. Consequently, we propose leveraging token embeddings from the middle layers, which have been shown to be contextual Phukan et al. (2024) and effective in representing concepts Wendler et al. (2024).
@@
  • We extend the method by introducing ContextualLens, which replaces the use of the logit lens with contextual embeddings from the middle layers of the VLM, leading to successful hallucination detection in categories that previously performed near random and improving detection accuracy in other categories.
@@ 2.2 Attribution and Grounded Visual Question Answering
ContextualLens is inspired by the training-free object hallucination detection and mitigation method proposed by Jiang et al. (2024), but with a crucial distinction. We utilize output token embeddings from intermediate layers instead of employing the logit lens probe. This modification enables us to overcome several limitations of the previous work, particularly by enhancing the representation of multi-token concepts and contextual relationships (§4.2).
@@ 4.2.1 Hallucination Detection
Given the answer tokens generated by the model, we compute the average embedding at a specific layer $l_{T}$, denoted as $h_{\text{Ans},l_{T}}$. Formally, let $\{t_{1},t_{2},\ldots,t_{k}\}$ be the sequence of answer tokens, and $h_{l_{T}}(t_{i})$ be the embedding of token $t_{i}$ at layer $l_{T}$. The average embedding $h_{\text{Ans},l_{T}}$ is computed as:
$h_{\text{Ans},l_{T}}=\frac{1}{k}\sum_{i=1}^{k}h_{l_{T}}(t_{i})$
Next, we evaluate the similarity between the answer embedding $h_{\text{Ans},l_{T}}$ and each image patch embedding from a specific layer $l_{I}$. Let $\{p_{1},p_{2},\ldots,p_{n}\}$ denote the set of image patches, and $h_{l_{I}}(p_{j})$ be the embedding of patch $p_{j}$ at layer $l_{I}$. We compute the score for each patch $p_{j}$ as the cosine similarity:
$\text{Score}(p_{j})=\text{CosineSim}(h_{\text{Ans},l_{T}},h_{l_{I}}(p_{j}))$
@@ 4.2.2 Grounded Visual Question Answering
Random LL Out Probs CL (Ours) Random LL Out Probs CL (Ours)
@@
Table 1: Comparison of mAP scores across different hallucination categories in the HQH dataset. The best-performing method in each category is bolded if it is significantly better than the second-best, and underlined if the difference is marginal.
@@ 6.1 Hallucination Detection
The performance of ContextualLens4.2.1) and the baseline methods (§5.3.1) on the HQH dataset is summarized in Table 1. Each category comprises 500 image-question pairs. To select the best layer combination for our method, we created a test and validation split with 250 examples each. The scores in Table 1 are reported for the test split. Further details on how to perform layer selection in real-world settings, including a robustness analysis using adversarial validation, are provided in Appendix C.
@@
Our method, which employs contextual embeddings, fares significantly better in these categories compared to the logit lens approach and often outperforms even output probabilities, sometimes by a substantial margin. We refer readers to Figures 6 and 7 for some qualitative examples.
For categories such as Action and OCR, where grounding actions (e.g., "Looking at the Television") or text (e.g., "Melrose Ave.") is necessary, the logit lens technique performs better than random but still falls short compared to output probabilities. This is likely due to the logit lens’s poor handling of multi-token objects. While objects which are usually single token may align relatively well, the complexity of representing actions or texts across multiple tokens poses challenges. Our approach, leveraging middle-layer embeddings known to better represent multi-token concepts Wendler et al. (2024), consistently outperforms the logit lens and, in many instances, surpasses output probabilities. See Figure 8 for a qualitative example.
Interestingly, output probabilities excel in the Count category compared to both ContextualLens and the logit lens. This category involves questions about the number of instances of a specific object. The inherent challenge for our approach and the logit lens is that visual elements often have low semantic overlap with numerical tokens (e.g., a "car" in the image is unlikely to semantically match "two" in the text). Consequently, output probabilities provide more accurate hallucination detection in such scenarios. A failure case is illustrated in Figure 9.
@@ 6.2.1 Qualitative Examples
Some qualitative results of our bounding box visual grounding technique are showcased in Figure 3. For instance, in Figure 3(a), we ground the phrase "Big Ben" to its corresponding location in the image. Our method effectively highlights the relevant portion by leveraging the latent conceptual representation of "Big Ben" intrinsic to the LMM, facilitating the mapping between textual descriptions and visual concepts. Another example is depicted in Figure 3(b). Remarkably, our method proficiently grounds answers within charts and infographics, a novel capability within the community to the best of our knowledge, achievable in a training-free manner. Additional qualitative examples are available in Appendix D.
@@ 7 Discussion & Conclusion
In this paper, we presented ContextualLens, an approach for detecting hallucinations in LMMs by leveraging contextual token embeddings from intermediate layers. We identified limitations in the state-of-the-art training-free technique, the logit lens, notably its poor handling of complex visual hallucinations involving attributes, comparisons, and spatial relations due to its reliance on non-contextual embeddings. Our proposed method employs contextual embeddings known to capture richer semantic information, improving hallucination detection across diverse categories.
Our experimental results on the HQH benchmark showed that ContextualLens enables detection in categories that previously performed near random and improves detection for multi-token concept representation. Furthermore, by introducing a novel grounding technique with highly precise bounding boxes, we advanced Zero-Shot Object Segmentation to a more general GVQA task, validating our approach’s effectiveness on TextVQA-X and VizWiz-G datasets.
@@ 8 Limitations & Future Work
Firstly, the validation of both hallucination detection and grounding has been conducted primarily on factual short answer VQA datasets. Extending this to more diverse and complex datasets remains an area for exploration. Secondly, ContextualLens is outperformed by output probabilities in the Count category. Future work could explore extensions to the method for abstractive scenarios such as counting. Lastly, although our method excels at highlighting precise evidence regions sufficient for human verification, improving recall remains an open challenge. Subsequent research could investigate leveraging our scoring mechanisms as priors for pre-trained segmentation models or improving the evaluation metrics for grounded visual question answering.
@@ B.2 Grounded Visual Question Answering
Appendix C Optimal Layer Selection
We use the hallucination detection task to discuss optimal layer selection, as its diverse categories make it ideal for evaluating robustness. The results in Table 1 were obtained using category-specific validation sets (250 examples per category). To test robustness, we conducted experiments with adversarial validation sets, where the category being tested was excluded from the validation data (1,500 examples: 250 × 6 categories). For InternLM-VL, we observed that Image layer embedding 13 and Text layer embedding 27 consistently ranked among the top two combinations across all categories. These results, summarized in Table 2, demonstrate that for all categories—except OCR—the performance closely matches that achieved with task-specific validation, indicating the robustness of layer selection. In the absence of task-specific data, a robust method for selecting optimal layers would be to identify those that rank highest during adversarial validation for available categories.
Category Task Specific Adversarial
Action 0.796 0.792
Attribute 0.825 0.833
Comparison 0.623 0.638
Count 0.885 0.896
Environment 0.811 0.813
Relation 0.755 0.752
OCR 0.772 0.744
Table 2: InternLM-VL mAP scores on HQH dataset for task-specific and adversarial validation.
Appendix D Additional Qualitative examples for Grounded Visual Question Answering
Figure 6: Top 20 image patches while detecting hallucination in attribute category using ContextualLens. We see that text tokens "Brown" are contextualized by the objects they refer to.
Figure 7: Top 5 image patches while detecting hallucination in comparison category using ContextualLens. We see that image patches corresponding to the younger boy’s hand are also highlighted in the second case.
Figure 8: Top 100 image patches while detecting hallucination in OCR category using ContextualLens. We see that image patches corresponding to multi-token "Michigan" are highlighted in the first case.
Figure 9: Top 20 image patches while detecting hallucination in count category using ContextualLens. We see that image patches highlighted in both cases are similar and the score for the second case is higher (0.891 < 0.892).